Show that w=z+e^z is analytic and hence find dw/dz
1
Expert's answer
2020-05-11T19:23:10-0400

Let's rewrite the function

"w(x+iy)=x+iy+e^{x+iy}=x+iy+e^x(\cos y+i\sin y)"

"u=Re(w)=x+e^x\cos y"

"v=Im(w)=y+e^x\sin y"

"\frac{\partial u}{\partial x}=1+e^x\cos y=\frac{\partial v}{\partial y}"

"\frac{\partial u}{\partial y}=-e^x\sin y=-\frac{\partial v}{\partial x}"

Therefore the function is analytic  and

"w'=1+e^z"